OpenStudy (anonymous):

How many solutions are there to the following system of equations? 6x – 9y = –9 y=2/3x+1 A-0 B-1 C-2 D-infinitely many

OpenStudy (141823):

I got the answer D Using substitution, I know y = 2/3x + 1 So I entered it into the equation 6x -9 (2/3x + 1) = -9 6x - 6x - 9 = -9 6x - 6x = 0 So, -9 = -9 Which is a true equation. Therefore, there are infinitely many solutions

OpenStudy (anonymous):

thank you, can you help me with another one